### v1

# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

Abstract
Neural network-based policies have demonstrated success in many robotic applications, but often lack human-explanability, which poses challenges in safety-critical deployments. To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification to describe a robot policy in a interpretable form. Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which do not give meaningful insights into the underlying policy. We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning. We also introduce three novel explainability evaluation metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification metrics. Our method is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing classification accuracy. This work bridges neural network policy learning with formal methods, contributing to safer and more transparent decision-making in robotics.
I Introduction
Recent innovations in learning-based methods, particularly those using neural networks, have significantly improved both high-level decision-making and low-level control in robotics. However, these neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount. To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2, 3], including post-hoc analysis techniques and methods to learn policies that are inherently explainable.
In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy. While existing approaches, like saliency maps [4] and feature attribution [5], provide valuable information, they often lack language-based explanations, which limits human-interpretability. Transformer-based language models have recently shown promising results for providing language-based explanations [6, 7, 8], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [9].
Formal methods offer an alternative that emphasizes verifiability. In particular, temporal logic has been used to specify robotic tasks and system constraints in an interpretable yet mathematically precise manner that supports verification [10, 11, 12]. Recent efforts have also used temporal logic, primarily linear temporal logic (LTL) and signal temporal logic (STL), to infer explanations of given trajectory data. For example, [13, 14, 15] use a decision tree classifier to infer the LTL or STL specification that best explains the difference between two classes of trajectories. [16] uses a greedy explanation search method over explanation candidates in LTL driven by policy similarity metrics. However, these methods are limited in their ability to infer specifications where certain predicates are more critical than others. For instance, [15] can infer such relations only for conjunctions. Weighted signal temporal logic (wSTL) [17] extends LTL and STL by introducing importance weights over predicates, which can highlight the relative priority of various task components. The expressiveness of wSTL thus makes it particularly suitable for describing robotic behaviors in settings with competing or hierarchical objectives.
Figure 1: Existing methods for neuro-symbolic post-hoc explainability produce specifications that are verbose (i.e., include unnecessary predicates), inconsistent across training seeds, and loose (e.g., “catch-all” specifications that do not provide useful insights). Our proposed network, TLNet, addresses these concerns.
Figure 2: We propose a neuro-symbolic approach for generating an explanation of a given robot policy using wSTL. Our network architecture, TLNet, includes a simplification process that ensures resulting explanations balance classification accuracy with conciseness, consistency, and strictness.
The presence of real-valued weights in wSTL has enabled the use of neural network approaches for inferring wSTL specifications that explain given trajectory data. These methods have shown promising results, where they show high classification performance for various applications, including anomaly detection in naval surveillance [18], fault diagnosis [19], and weather forecasting [20]. However, existing work only focuses on classification accuracy, which does not fully capture the actual explanability of learned specifications. For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications that do not fully capture the task or constraints). Some works address verbosity by incorporating heuristic simplification methods, like only keeping the top-$k$ weighted predicates [18, 20] or using a greedy process to remove weights [21]. However, as we show in our empirical results, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
In this work, we propose TLNet, a neural network architecture for inferring wSTL explanations of robot policies that are concise (i.e., include as few predicates as needed), consistent (i.e., produce similar explanations across training seeds), and strict (i.e., contain as few disjunctions as needed). Our method targets a subclass of wSTL structures with arbitrary conjunctions and disjunctions of predicates in temporal clauses. TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained, as shown in Figure 2. This process includes a predicate filtering step, a regularized loss function, and iterative weight pruning. We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduce learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices. We demonstrate our method in three simulated robotics environments that vary in continuity, dimensionality, task structure, and policy complexity, and show that it generally outperforms baselines with respect to classification accuracy and our proposed explainability metrics. TLNet bridges the gap between neural network-based policy learning and providing formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
II Preliminaries
Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},..,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon. We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
II-A Weighted Signal Temporal Logic
We consider wSTL syntax [17] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a real-valued function $f:S\rightarrow\mathbb{R}$ over $S$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively. For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator. For temporal operators, we define the weights as $w:=[w_{k}]_{k\in I}\in\mathbb{R}^{\lvert I\rvert}_{>0}$ for interval $I$. For notational simplicity, we use the following notation interchangeably: $\bigwedge_{i=1}^{N}{}^{w}\phi_{i}=\bigwedge_{i=1}^{N}w_{i}\phi_{i}$ and $\bigvee_{i=1}^{N}{}^{w}\phi_{i}=\bigvee_{i=1}^{N}w_{i}\phi_{i}$.
II-B Normalized wSTL
Our neural network architecture proposed in Section III-A requires the weights within each operator to be normalized. We discuss the reasons for this requirement in Section III-E. Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=1$ and $\sum_{k\in I}w_{k}=1$, where $0\leq w_{i},w_{k}\leq 1$. We allow the weights to be zero because the zero weights do not influence the overall robustness [18].
To avoid large input values to our neural network, we also define normalized quantitative satisfaction semantics, or robustness, by extending [17] as follows,
$r(\psi,\tau,t)$ ::=:= $\begin{cases}\frac{f(s_{t})-c}{\sup_{s\in S}f(s)-c}&\text{if $f(s_{t})-c\geq 0$,}\\ \frac{f(s_{t})-c}{\inf_{s\in S}f(s_{t})-c}&\text{otherwise}\end{cases},$ (1)
$r(\neg\phi,\tau,t)$ ::=:= $-r(\phi,\tau,t),$ (2)
$r\left(\bigwedge_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\otimes^{\wedge}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (3)
$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}).$ (4)
$r\left(\mathcal{G}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\otimes^{\mathcal{G}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (5)
$r\left(\mathcal{F}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\oplus^{\mathcal{F}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (6)
where the aggregation functions $\otimes^{\wedge}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\vee}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\mathcal{G}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$, and $\otimes^{\mathcal{F}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$ correspond to the operators $\wedge,\vee,\mathcal{G}$, and $\mathcal{F}$, respectively. For notational simplicity, we interchangeably use $r(\phi,\tau,t)$ and $r_{\tau t}^{\phi}$.
Building from [18], we define the aggregation functions as follows,
$\otimes^{\wedge}(w,[r_{\tau t}^{\phi_{i}}]_{i=1:N})$ $:=\frac{\sum_{i=1}^{N}w_{i}r_{\tau t}^{\phi_{i}}\exp(\frac{-r_{\tau t}^{\phi_{i}}}{\sigma})}{\sum_{i=1}^{N}w_{i}\exp(\frac{-r_{\tau t}^{\phi_{i}}}{\sigma})},$ (7)
$\oplus^{\vee}(w,[r_{\tau t}^{\phi_{i}}]_{i=1:N})$ $:=\frac{\sum_{i=1}^{N}w_{i}r_{\tau t}^{\phi_{i}}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})}{\sum_{i=1}^{N}w_{i}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})},$ (8)
$\otimes^{\mathcal{G}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
$\oplus^{\mathcal{F}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (10)
where $\sigma>0$ is a scaling constant.
II-C Problem Statement
We consider the following binary classification task. Let $\mathcal{T}=\Set{(\tau_{i},l_{i})}{i=1,\dots,N_{\mathrm{data}}}$ be a labeled dataset, where $\tau_{i}\in\mathbb{R}^{n\times H}$ is the $i^{th}$ trajectory with label $l_{i}\in\{1,-1\}$ and $N_{\mathrm{data}}\in\mathbb{Z}_{>0}$ is the number of data points in the dataset.
Problem 1.
Given a dataset $\mathcal{T}$, find a wSTL specification $\phi$ that balances classification accuracy with conciseness, consistency, and strictness.
III Methodology
We first introduce our neural network architecture, TLNet, which allows us to represent wSTL structures in a fully differentiable form. We then introduce our simplification process, which enhances the explainability of the generated explanations, and our trajectory sampling process. Finally, we propose novel evaluation metrics for explanability—conciseness, consistency, and strictness—and provide theoretical justification for our network architecture. Algorithm 1 summarizes our overall approach.
Algorithm 1 Explanation Generation
1:Dataset $\mathcal{T}$, predicates $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$, similarity threshold $S_{\mathrm{th}}$, number of pruning iterations $N_{\mathrm{pr}}$, number of weights to prune $N_{\mathbf{w}}$, loss function $\tilde{\mathcal{L}}.$
2:Explanation $\phi$
3:Evaluate robustness of trajectories in $\mathcal{T}$ using $\Psi$.
4:Filter out predicates whose similarities are above $S_{\mathrm{th}}$.
5:Initialize the neural network $f_{\mathbf{w}}$ with parameters $\mathbf{w}$.
6:Optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
7:for $i=1:N_{\mathrm{pr}}$ do
8:     Prune zeroed and smallest $N_{\mathbf{w}}$ parameters in $\mathbf{w}$.
9:     Re-normalize $\mathbf{w}$.
10:     Optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
11:end for
12:return $\phi$ generated from $\mathbf{w}$ and $\Psi$ using Equation 11.
III-A Neural Network Architecture
Let $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a predefined set of atomic predicates and $\Psi_{\neg}=\Set{\neg\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a set of their negations. We focus on representing wSTL structures composed of temporal operators containing these predicates in conjunctive normal form (CNF), as follows,
$\phi=0.5\mathcal{F}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]\\ \wedge 0.5\mathcal{G}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}}\psi_{j}\right],$ (11)
where $w_{ij}^{\mathcal{F}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates defining specification $\phi$. We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [22, 16]. We define the weights $w^{U}$ for temporal operators over interval $I=[0,H]$ to be $w^{U}=[w^{U}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$. We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation.
The architecture of TLNet is designed to ensure the built-in architectural interpretability of the network weights, achieved by the bijective correspondence of the network weights and the importance weights of Equation 11. In Section III-E, we detail the theoretical justification of the formula manipulation to aggregate the weights of the conjunctive and disjunctive clauses in Equation 11. The output of TLNet is the robustness of the input trajectory with respect to the specification defined by the current network weights.
An example TLNet architecture for a dataset containing two predicates is shown in Figure 3. The specification represented by this architecture is $\phi=0.5\mathcal{F}_{I}^{w^{U}}[(w_{11}^{\mathcal{F}}\psi_{1}\vee w_{12}^{\mathcal{F}}\psi_{2}\vee w_{13}^{\mathcal{F}}\neg\psi_{1}\vee w_{14}^{\mathcal{F}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{F}}\psi_{1}\vee w_{22}^{\mathcal{F}}\psi_{2}\vee w_{23}^{\mathcal{F}}\neg\psi_{1}\vee w_{24}^{\mathcal{F}}\neg\psi_{2})]\wedge 0.5\mathcal{G}_{I}^{w^{U}}[(w_{11}^{\mathcal{G}}\psi_{1}\vee w_{12}^{\mathcal{G}}\psi_{2}\vee w_{13}^{\mathcal{G}}\neg\psi_{1}\vee w_{14}^{\mathcal{G}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{G}}\psi_{1}\vee w_{22}^{\mathcal{G}}\psi_{2}\vee w_{23}^{\mathcal{G}}\neg\psi_{1}\vee w_{24}^{\mathcal{G}}\neg\psi_{2})]$.
Figure 3: Example architecture of the proposed neural network when there are two predicates to consider ($\psi_{1}$ and $\psi_{2}$). $\mathbf{w}^{\mathcal{F}}$ and $\mathbf{w}^{\mathcal{G}}$ are weight vectors that are optimized. $\phi$ is the inferred explanation to the given trajectory.
III-B Simplification Process
A key component of TLNet is its simplification process, which aims to produce more human-interpretable wSTL explanations while preserving classification performance. This process consists of three steps: a predicate filter, loss regularization, and weight pruning.
The predicate filter eliminates predicates that exhibit similar behavior across both positive and negative trajectory sets, as they contribute little to discriminating between those sets and may introduce unnecessary complexity. Given a dataset of trajectories $\mathcal{D}$ and a predicate $\psi\in\Psi$, we define the robustness distribution as
$d(\mathcal{D},\psi)=\frac{1}{|\mathcal{D}|}\left[\lvert\mathcal{D}_{1}\rvert,\lvert\mathcal{D}_{2}\rvert,\lvert\mathcal{D}_{3}\rvert\right],$ (12)
where
$\mathcal{D}_{1}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})\geq 0,\forall t^{\prime}\in t+I},$ (13)
$\mathcal{D}_{2}$ $:=\mathcal{D}\setminus(\mathcal{D}_{1}\cup\mathcal{D}_{3}),$ (14)
$\mathcal{D}_{3}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})<0,\forall t^{\prime}\in t+I}.$ (15)
Let $\mathcal{T}_{p}=\{(\tau_{i},l_{i})|l_{i}=1\}_{i=1:N_{\mathrm{data}}}$ be the positive set of trajectories and $\mathcal{T}_{n}=T\setminus T_{p}$ be the negative set. For each predicate $\psi\in\Psi$, we then calculate its robustness distributions for $\mathcal{T}_{p}$ and $\mathcal{T}_{n}$$d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, respectively—and their cosine similarity $S:=\frac{d(\mathcal{T}_{p},\psi)\cdot d(\mathcal{T}_{n},\psi)}{\lvert d(\mathcal{T}_{p},\psi)\rvert\lvert d(\mathcal{T}_{n},\psi)\rvert}$. As a result, predicates (and their negation) with a similarity score exceeding a user-defined threshold $S_{\mathrm{th}}$ are removed from $\Psi$ (and $\Psi_{\neg}$) prior to optimization.
[…]
$R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (16)
The intuition behind the max operation is to choose the larger sum of weights of a predicate and those of its negation. Next, we define disjunctive clause regularizer $R_{\mathrm{D}}$. Given a temporal clause with importance weights $\mathbf{w}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$, $R_{\mathrm{D}}$ is defined as,
$R_{\mathrm{D}}(\mathbf{w}):=\sum_{i=1}^{N_{\mathrm{AP}}-1}\sum_{j=i+1}^{N_{\mathrm{AP}}}\sum_{k=1}^{2N_{\mathrm{AP}}}w_{ik}w_{jk},$ (17)
where $w_{ij},w_{jk}\in\mathbf{w}$. The intuition behind the formulation is to differentiate the used predicates in a disjunctive clause in CNF from the predicates in the other disjunctive clauses within a temporal clauses.
Putting these regularizers together, the overall loss $\tilde{\mathcal{L}}$ is
~$\tilde{\mathcal{L}}(\mathcal{L},\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})=\mathcal{L}+\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})\\ +\lambda_{R_{\mathrm{D}}}[R_{\mathrm{D}}(\mathbf{w}^{\mathcal{F}})+R_{\mathrm{D}}(\mathbf{w}^{\mathcal{G}})],$ (18)
where regularizer coefficients $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are small real values, and $\mathcal{L}$ is a desired loss function. In this work, we use the loss $\mathcal{L}$ proposed by [18],
$\mathcal{L}(\phi)=\sum_{j=1}^{N_{\mathrm{data}}}\exp(-\zeta l_{j}r(\phi,\tau_{j},t))$ (19)
The weight pruning stage further simplifies the currently inferred wSTL specification by reducing the number of active predicates. This step proceeds in two phases. First, all weights with a value of zero are pruned, ensuring that these predicates are excluded from subsequent optimization steps. Second, among the remaining nonzero weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified parameter—are removed to eliminate the least contributing parts of the specification. After pruning, the remaining nonzero weights are normalized such that their sum equals one.
III-C Trajectory Sampling
As described in Section II, the neural network is trained as a binary classifier to distinguish between positive and negative trajectories. Positive trajectories are obtained by executing the target policy in the environment of interest. Alternatively, they can be sampled from existing datasets, depending on the application setting. Negative trajectories are generated using a random-walk policy in this work, serving as contrasting samples for the classifier. However, the choice of policy used for generating negative trajectories is flexible and can be adapted based on the user’s objectives or the specific characteristics of the environment.
III-D Explanability Evaluation Metrics
Beyond classification performance, TLNet emphasizes the human-interpretability and alignment of inferred wSTL explanations with underlying policy behavior. While alignment is ultimately subjective and often requires qualitative assessment, we introduce three quantitative metrics—conciseness, consistency, and strictness—to characterize key aspects of explanation quality.
Let $\phi$ be an wSTL explanation we want to evaluate and $N$ be the number of temporal clauses that we expect to appear in the explanation. For example, our specifications expect $N=2$ temporal clauses. First, we define a conciseness metric as,
$\mathrm{Conciseness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{m_{n}}&\text{if $m_{n}\neq 0$},\\ 0&\text{otherwise},\end{cases}$
where $m_{n}$ is the number of predicates present in $n$-th temporal clause of $\phi$. Conciseness ranges from 0 to 1, with higher scores implying a shorter explanation with at least one predicate inside each expected temporal clause.
We now define a consistency metric. Let $\Phi=\{\phi_{k}\}_{k=1:K}$ be a set of $K$ explanations for the same policy (e.g., from different training seeds), and $\phi_{kn}$ be the $n$-th temporal clause ($n\in[1,N]$) expected to appear within $\phi_{k}$. We define a set of unique temporal clause structures $\Xi^{n}$ for a set of the $n$-th temporal clauses $\Set{\phi_{kn}}_{k=1:K}$. Each unique structure $\xi^{n}\in\Xi^{n}$ has $c_{\xi^{n}}$ temporal clauses satisfying $\xi^{n}$. Then, consistency of $\Phi$ is defined as
$\mathrm{Consistency}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\xi^{n}\in\Xi^{n}}\frac{c_{\xi^{n}}}{\lvert\Xi^{n}\rvert}&\text{if $\Xi^{n}\neq\emptyset$},\\ 0&\text{otherwise}.\end{cases}$
Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
Finally, we introduce a strictness metric. Let $\phi_{n}$ be the $n$-th temporal clause of an explanation $\phi$ and $P$ be the number of potentially present predicates in $\phi$ (e.g., $P=2N_{\mathrm{AP}}$ in TLNet). Denote the numbers of conjunctions and disjunctions in $\phi_{n}$ as $C_{n}$ and $D_{n}$, respectively. Then, strictness is defined as,
$\mathrm{Strictness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{P-C_{n}+D_{n}}&\text{if $\phi_{n}$ is not empty},\\ 0&\text{otherwise.}\end{cases}$
Strictness ranges from 0 to 1, with higher scores implying stricter logical expressions by rewarding conjunctions and penalizing disjunctions.
III-E Weight Distribution in wSTL
A key assumption in our neural network architecture is that weights can be distributed across Boolean and temporal clauses. We now prove that such distribution is theoretically sound. Recall that an arbitrary Boolean formula $\phi$ can be expressed in CNF as,
$\phi=\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}(\neg)\psi_{j},$ (20)
where $n,m_{i}\in\mathbb{Z}_{>0}$ and $\psi_{ij}$ are predicates. The normalized weights of an arbitrary CNF wSTL formula can be syntactically distributed across Boolean operators as follows.
Theorem 1 (Syntactic Weight Distributivity of Boolean Operators).
A normalized wSTL formula $\phi$ in CNF can be represented as follows,
$\phi=\bigwedge_{i=1}^{n}w_{i}\bigvee_{j=1}^{m_{i}}w_{ij}(\neg)\psi_{ij}\equiv\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}w_{ij}^{\prime}(\neg)\psi_{ij},$ (21)
where $n,m_{i}\in\mathbb{Z}_{>0}$, $w_{i},w_{ij}\in\mathbb{R}_{>0}$, $w_{ij}^{\prime}:=w_{i}w_{ij}$, $\sum_{i=1}^{n}w_{i}=\sum_{j=1}^{m_{i}}w_{ij}=1$, and $\psi_{ij}$ are predicates. Furthermore, the weights $w_{ij}^{\prime}$ satisfy $\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}w_{ij}^{\prime}=1$.
Proof.
[…]
$\wedge w_{2}(w_{21}(\neg)\psi_{21}\vee...\vee w_{2m_{2}}(\neg)\psi_{2m_{2}})$
... $\wedge w_{n}(w_{n1}(\neg)\psi_{n1}\vee...\vee w_{nm_{n}}(\neg)\psi_{nm_{n}})$
$=(w_{1}w_{11}(\neg)\psi_{11}\vee...\vee w_{1}w_{1m_{1}}(\neg)\psi_{1m_{1}})$
$\wedge(w_{2}w_{21}(\neg)\psi_{21}\vee...\vee w_{2}w_{2m_{2}}(\neg)\psi_{2m_{2}})$
... $\wedge(w_{n}w_{n1}(\neg)\psi_{n1}\vee...\vee w_{n}w_{nm_{n}}(\neg)\psi_{nm_{n}})$
$=\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}w_{i}w_{ij}(\neg)\psi_{j}=\{\text{RHS}\}.$
IV Results
We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks. We consider three baselines: {outline} \1 A greedy pruning strategy proposed by [21], which independently removes individual network weights that do not affect classification accuracy post-optimization. \1 Top-$k$ pruning strategies (top-3 and top-5) used by [23, 20], which only retain the $k$ largest network weights. All baselines were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison. For each policy we aimed to explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy. Each dataset was split 80% and 20% for training and evaluation, respectively.
(a) CtF environment.
(b) Obstructed fetch push.
(c) Chased robot navigation.
Figure 4: Screenshots from our environments.
IV-A Test Environments
We consider the three environments shown in Figure 4, which include discrete and continuous state-action spaces, navigation and manipulation tasks, and varied numbers of predicates.
IV-A1 Capture-the-flag (CtF)
CtF is a discrete grid-world environment with adversarial dynamics where two agents compete to capture each other’s flags, based on [16]. If both agents occupy adjacent cells in a given territory, the defending agent eliminates the other with 75% probability. We considered blue agent policies optimized for five scenarios. The first four considered different red agent heuristic policies (capture, fight, patrol, roomba) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$). The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0). The red agent policies are defined as follows: capture aggressively targets the flag; fight pursues the blue agent; patrol defends a boundary region; and roomba switches between capturing and defending. In all red agent policies, the red agent performs random actions 25% of the time. We defined eight atomic predicates based on euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”) and binary defeat (“df”) status $\{1,-1\}$ of the red agent: $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.
IV-A2 Fetch Push
We also consider a continuous control task based on the Fetch mobile manipulator [24]. The robot must push a blue block (“b”) to a red target (“t”) using its gripper (“g”), without disturbing a black obstacle (“o”). The state and action spaces are continuous ($\mathcal{S}\subset\mathbb{R}^{55},a\in\mathbb{R}^{4}$). Eight predicates were defined based on the distances between the gripper, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
IV-A3 Robot Navigation
We extended the CarGoal-v0 environment [25] to create a more complex adversarial navigation task. The ego robot (“e”) must reach a green goal (“g”) while avoiding a violet chaser (“c”) moving with a fixed speed ($v_{c}=7.5\times 10^{-4}$). Static hazards (“h”) and movable vase (“v”) are present but not penalized. The state and action dimensions are $s\in\mathbb{R}^{88}$ and $a\in\mathbb{R}^{2}$, respectively. Ten distance-based predicates were defined, involving euclidean distances between the robot, the chaser, the goal and the hazards: $\psi_{\mathrm{ec}}$, $\psi_{\mathrm{eg}}$, $\psi_{\mathrm{eh}}$, $\psi_{\mathrm{ev}}$, $\psi_{\mathrm{gc}}$, $\psi_{\mathrm{gh}}$, $\psi_{\mathrm{gv}}$, $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$, $\psi_{\mathrm{vc}}$.
To test robustness to noise, we included predicates irrelevant to task success (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) and evaluated whether TLNet could eliminate them in its explanations.
IV-B Policy Optimization
TLNet is a general framework to generate explanations of robot policies, though, we focused on policies optimized by deep reinforcement learning algorithms in this work. We used Stable-Baselines3 [26] for training all policies. For the CtF and navigation environments, PPO [27] was used; for the Fetch push task, we used SAC [28] with Hindsight Experience Replay (HER) [29]. Hyperparameters were taken from prior work [16, 30, 25]. Blue agent policies in CtF achieved over 97% success, except the roomba case, which reached 77.5% success in 1,000 evaluation episodes. For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explanability metrics.
TABLE I: Baseline Comparison of Representative Generated Explanations.
Scenarios TLNet (Ours) Greedy Top-3 Top-5
CtF Capture
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
$(0.09\neg\psi_{\mathrm{ba,bt}}\vee 0.07\neg\psi_{\mathrm{ra,bt}})]\wedge$
$0.5\mathcal{G}[0.36\psi_{\mathrm{ba,rf}}\vee 0.64\neg\psi_{\mathrm{ra,bf}}]$
$\mathcal{F}[0.73\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.27\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
CtF Capture 0
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[0.08\psi_{\mathrm{ba,rf}}\wedge(0.40\psi_{\mathrm{ba,rf}}\vee 0.31\neg\psi_{\mathrm{ra,bt}})\wedge$
$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}[(0.33\psi_{\mathrm{ba,rf}}$
$\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
$\mathcal{F}[0.67\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
CtF Fight
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
$0.5\mathcal{F}[0.77\psi_{\mathrm{ba,rf}}\wedge(0.15\psi_{\mathrm{ba,rf}}\vee 0.08\psi_{\mathrm{ra,bf}})]\wedge$
$0.5\mathcal{G}[(0.15\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})\wedge 0.06\psi_{\mathrm{ba,rf}}$
$\vee 0.04\neg\psi_{\mathrm{ra,df}})\wedge(0.06\psi_{\mathrm{ba,rf}}\vee 0.06\neg\psi_{\mathrm{ba,bt}}\vee$
$0.04\neg\psi_{\mathrm{ra,df}})\wedge(0.02\psi_{\mathrm{ba,ra}}\vee 0.19\psi_{\mathrm{ba,rf}}\vee$
$0.16\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})]$
$\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$
CtF Patrol
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
$0.5\mathcal{F}[0.35\psi_{\mathrm{ba,rf}}\wedge(0.09\psi_{\mathrm{ba,rf}}\vee 0.55\psi_{\mathrm{ra,df}})]\wedge$
$0.5\mathcal{G}[(0.29\psi_{\mathrm{ba,rf}}\vee 0.04\psi_{\mathrm{ra,bf}}\vee 0.07\neg\psi_{\mathrm{ra,df}})\wedge$
$(0.30\psi_{\mathrm{ba,rf}}\vee 0.19\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ba,bt}})\wedge$
$0.04\psi_{\mathrm{ra,bf}}\wedge 0.05\neg\psi_{\mathrm{ra,bf}}]$
$\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$
CtF Roomba
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.35\psi_{\mathrm{ba,rf}}\vee 0.65\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
$\wedge(0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.05\neg\psi_{\mathrm{ra,df}})\wedge(0.18\psi_{\mathrm{ba,rf}}\vee$
$0.03\psi_{\mathrm{ra,df}}\vee 0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ra,bt}})]$
$\mathcal{F}[1.0\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.48\neg\psi_{\mathrm{ba,bt}}$
$\wedge 0.52\psi_{\mathrm{ba,rf}}]$
Fetch Push
$0.5\mathcal{F}[1.0\psi_{\mathrm{bt}}]\wedge$
$0.5\mathcal{G}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
$0.5\mathcal{F}[0.74\psi_{\mathrm{bt}}\wedge(0.10\psi_{\mathrm{bt}}\vee 0.16\psi_{\mathrm{od}})]\wedge$
$0.5\mathcal{G}[1.0\neg\psi_{\mathrm{bd}}]$
$\mathcal{G}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}[1.0\psi_{\mathrm{bt}}]$
Robot Navi. $0.5\mathcal{F}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$
TABLE II: Baseline Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics.
Conciseness Consistency Strictness Accuracy
Scenarios Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5
CtF Capture 0.41 0.21 0.20 0.22 0.33 0.08 0.08 0.13 0.33 0.08 0.08 0.13 0.95 0.95 0.95 0.94
CtF Capture 0 0.42 0.21 0.20 0.30 0.45 0.11 0.15 0.24 0.45 0.11 0.15 0.24 0.95 0.95 0.95 0.95
CtF Fight 0.39 0.18 0.25 0.25 0.68 0.09 0.50 0.50 0.68 0.09 0.50 0.50 0.94 0.92 0.94 0.93
CtF Patrol 0.38 0.18 0.25 0.28 0.63 0.05 0.50 0.53 0.63 0.05 0.50 0.53 0.95 0.93 0.94 0.94
CtF Roomba 0.39 0.12 0.21 0.16 0.27 0.06 0.10 0.09 0.27 0.06 0.10 0.09 0.87 0.89 0.88 0.88
Fetch Push 0.42 0.31 0.25 0.25 1.00 0.28 0.50 0.50 1.00 0.28 0.50 0.50 0.99 0.97 0.97 0.97
Robot Navi. 0.48 0.22 0.25 0.25 0.68 0.15 0.50 0.50 0.68 0.15 0.50 0.50 0.97 0.96 0.96 0.96
IV-C Experimental Results
Table I presents representative explanations for each scenario, comparing our method with three baseline. Across all scenarios, TLNet consistently inferred both task ($\mathcal{F}$) and constraint ($\mathcal{G}$) clauses, while the top-$k$ methods often captured only one of the two. This structural incompleteness stems from the fact that top-$k$ pruning lacks guarantees on preserving clause-level composition, despite yielding concise formulas with high classification accuracy.
Our method produced well-structured explanations across environments. For instance, task clauses included $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”) in CtF, and $\psi_{\mathrm{bt}}$ (“block reaches the target”) and $\psi_{\mathrm{eg}}$ (“robot reaches the target” in Fetch and robot navigation. Constraint clauses correctly included safety conditions like $\neg\psi_{\mathrm{ra,bf}}$ (“the red agent does not capture the blue flag”) or $\neg\psi_{\mathrm{bd}}$ (“the block is not dropped”), with the exception of the navigation domain, where structural guarantees were again lacking in the baselines.
As shown in Table II, our method outperformed all baselines in conciseness, consistency, and strictness, with only marginal differences in classification accuracy. On average, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness. The lower accuracy in the CtF Roomba scenario can be attributed to the suboptimality of the learned policy, which reduced the separability between positive and negative trajectories. These results support the qualitative findings above, validating the explainability of our method’s outputs.
Across CtF scenarios, TLNet consistently captured the task $\psi_{\mathrm{ba,rf}}$. The capture and capture 0 scenarios yielded identical constraint structures: $\mathcal{G}(0.30\,\psi_{\mathrm{ba,rf}}\vee 0.70\,\neg\psi_{\mathrm{ra,bf}})$ and $\mathcal{G}(0.31\,\psi_{\mathrm{ba,rf}}\vee 0.69\,\neg\psi_{\mathrm{ra,bf}})$, reflecting their similar optimized behaviors despite reward differences. In contrast, the fight scenario reflected the learned avoidance strategy through the constraint $\mathcal{G}(0.33\,\psi_{\mathrm{ba,rf}}\vee 0.67\,\neg\psi_{\mathrm{ba,ra}})$, emphasizing avoidance of battles.
The inferred explanation for Fetch Push was: “With importance 0.5, eventually the block reaches the target; with importance 0.5, globally, the grip contacts the block (0.41) or the block is on the target (0.59)”. While the task was correctly captured, the constraint clause was only partially valid due to the use of smoothed min/max activations. Since the grip reaches the block early in the episode, negative robustness signals were underemphasized. Additionally, the predicate $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”) was not inferred, likely due to minimal violations in both positive and random-walk trajectories. This illustrates a limitation of the binary classification formulation and the trajectory sampling strategy. Future work could address this by using adversarial or policy-guided sampling to generate more informative negative examples.
For the navigation task, TLNet inferred a clear and aligned explanation: “With importance 0.5, eventually the robot reaches the goal; with importance 0.5, globally, it avoids the chaser.” The method successfully ignored irrelevant predicates (e.g., $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$), demonstrating robustness against unnecessary input predicates and alignment with the tasks and constraints.
TABLE III: Ablation Study of CtF Capture Scenario.
Metrics Full No reg. No pruning No filter
Accuracy 0.951 0.954 0.776 0.947
Conciseness 0.408 0.396 0.147 0.289
Consistency 0.325 0.233 0.2 0.175
Strictness 0.117 0.117 0.063 0.106
Table III presents ablation results on the CtF Capture scenario. The full pipeline achieved the highest conciseness, consistency, and strictness. A variant without regularization slightly outperformed in accuracy, which is expected, as regularization biases optimization toward simpler explanations. However, the accuracy trade-off was minimal ($<0.01$), while consistency improved by 39.4% and conciseness by 3.0%. Removing the predicate filter significantly reduced conciseness (27.2%) and consistency (46.1%), highlighting its importance. The most impactful component was pruning, particularly in reducing explanation size—nearly tripling conciseness. Overall, the ablation study confirms that each step of our simplification process plays a vital role in producing interpretable, structurally sound explanations without compromising performance.
V Conclusion
We introduced a neuro-symbolic framework, TLNet, for generating weighted Signal Temporal Logic (wSTL) explanations of robot policies, addressing limitations of trajectory classification-based methods. Our approach directly explains policies while refining explanations through predicate filtering, loss regularization, and weight pruning. To assess quality beyond classification metrics, we proposed human-interpretability metrics—conciseness, consistency, and strictness. Experiments in three simulated robotics environments showed superior explanation explanability with high accuracy, with an ablation study confirming our simplification process’s effectiveness. By bridging neural policy learning and formal logic-based explanations, our work advances transparency in robotics, with future directions including more complex wSTL structures, human-in-the-loop refinement, and real-world applications.